# Kinetic temperature and radial flow velocity estimation using identified hadrons and light (anti-)nuclei produced in relativistic heavy-ion collisions at RHIC and LHC

### v2

@@ Abstract
We report the investigation of the kinetic freeze-out properties of identified hadrons ($\pi^{\pm}$, $K^{\pm}$ and $p(\bar{p})$) along with light (anti-)nuclei $d(\bar{d})$, $t(\bar{t})$ and ${}^{3}He$ in relativistic heavy-ion collisions at RHIC and LHC energies. A simultaneous fit is performed with the Blast-Wave (BW) model to the transverse momentum ($p_{\rm T}$) spectra of identified hadrons together with light (anti-)nuclei produced in Au+Au collisions at $\sqrt{\mathrm{s}_{{}_{\mathrm{NN}}}}$ = 7.7 – 200 GeV at the RHIC and in Pb+Pb collisions at $\sqrt{\mathrm{s}_{{}_{\mathrm{NN}}}}$ = 2.76 TeV at the LHC. The energy and centrality dependence of freeze-out parameters, i.e., kinetic freeze-out temperature ($T_{kin}$) and collective flow velocity $\langle\beta\rangle$ has been studied. It is observed that light (anti-)nuclei also participate in the collective expansion of the medium created in the collision when included in a common fit with the light hadrons. We observe a marginal rise in $T_{kin}$ and a slight decrease in $\langle\beta\rangle$ when compared to the values obtained from the fit to light hadrons. A similar $\langle\beta\rangle$ and significantly larger $T_{kin}$ is observed when a fit is performed to only protons and light (anti-)nuclei. Additionally, we also observe that the $T_{kin}$ increases from central to peripheral collisions, which is consistent with the argument of a short-lived fireball in peripheral collisions. Whereas the $\langle\beta\rangle$ shows a decreasing trend from central to peripheral collisions indicating a more rapid expansion in the central collisions. Both, $T_{kin}$ and $\langle\beta\rangle$ show a weak dependence on the collision energy at most energies.
@@ III.1 Transverse Momentum ($p_{\rm T}$) Spectra
For light hadrons, we repeated the analysis presented in Ref. [56] for a simultaneous fit to the $p_{\rm T}$ spectra with BW model. The obtained parameters from the BW fit are consistent with the previously reported results [56]. The low $p_{\rm T}$ part of the pion spectra is significantly influenced by resonance decays therefore, $p_{\rm T}$ $<0.5$ GeV/$c$ is excluded from the fit. Conversely, at higher $p_{\rm T}$, eventual hard contributions may become significant, which are not anticipated to be effectively described by the BW model [65]. Therefore, the sensitivity of the BW model parameters especially $T_{kin}$ is contingent upon the choice of $p_{\rm T}$ ranges used for fitting [57]. The fit range for light hadrons have been limited according to Ref. [56] (0.5 – 1.3 GeV/$c$, 0.25 – 1.4 GeV/$c$, 0.4 – 1.3 GeV/$c$ for $\pi^{\pm},K^{\pm}$ and $p(\bar{p})$ respectively). However, for light (anti-)nuclei the spectrum is fitted up to the maximum available $p_{\rm T}$ range. The fit ranges can indeed affect the extracted $T_{kin}$, especially at lower $p_{\rm T}$. On the other hand, the maximum $p_{\rm T}$ described by the BW model should increase from peripheral to central collisions, or increase from lower to higher collision energies. This is because the particles are expected to be more thermalized when the volume and lifetime are larger. This suggests that different fit ranges would be appropriate for peripheral and central collisions, or for different collision energies. More details are reported in Refs. [70, 71].
@@ III.2 Kinetic Freeze-out Parameters
Figures 3 and  4 show the centrality dependence of temperature ($T_{kin}$) and radial flow velocity ($\langle\beta\rangle$), respectively. Parameters are extracted by simultaneous fits to the $p_{\rm T}$ spectra of light hadrons along with light (anti-)nuclei in Au+Au collisions at $\sqrt{\mathrm{s}_{{}_{\mathrm{NN}}}}$ = 7.7 – 200 GeV from RHIC [56] and Pb+Pb collisions at $\sqrt{\mathrm{s}_{{}_{\mathrm{NN}}}}$ = 2.76 TeV [57] from LHC. For simplicity, the centrality classes of $\pi^{\pm}$, $K^{\pm}$, $p(\bar{p})$ are merged and refitted to make it consistent with those used for $d(\bar{d})$ and $t(\bar{t})$ measurements, i.e. 0–10%, 10–20%, 20–40% and 40–80%. Due to unavailability of the triton $p_{\rm T}$ spectra from Pb+Pb collisions at $\sqrt{\mathrm{s}_{{}_{\mathrm{NN}}}}$ = 2.76 TeV, we included ${}^{3}He$ in the fits instead. Blue markers represent the published results for $\pi^{\pm}$, $K^{\pm}$, and $p(\bar{p})$. Green markers depict $\pi^{\pm}$, $K^{\pm}$, and $p(\bar{p})$ data with combined centrality classes. Red markers correspond to $T_{kin}$ of $\pi^{\pm}$, $K^{\pm}$, $p(\bar{p})$, and $d(\bar{d})$. It is evident from Fig. 3 that $T_{kin}$ exhibits an increasing trend as a function of centrality for all collision energies. Furthermore, the inclusion of heavier particle species, such as light (anti-)nuclei, results in slightly higher temperatures compared to those of light hadrons. This effect is more prominent at energies $\sqrt{\mathrm{s}_{{}_{\mathrm{NN}}}}$geq 62.4 GeV. Fig. 4 shows that fitting both light hadrons with light (anti-)nuclei together leads to a smaller values of $\langle\beta\rangle$ as compared to those when only light hadrons are considered in the fits. To study the influence of light (anti-)nuclei data on the global characteristics, we repeat the fit by excluding $d(\bar{d})$ and $t$ simultaneously exhibiting similar trend with slight smaller value of $T_{kin}$ and sightly larger values of $\langle\beta\rangle$. This suggest that heavier particles decouple earlier in time from the fireball compared to the light hadrons due to higher masses and are more sensitive to the effects of the radial flow. Other effects, like elliptic flow ($v_{2}$) of deutrons can also be used study the distinction between coalescence and statistical production. More details are discussed in Ref. [78]. Moreover, the $T_{kin}$ in mid-central 20–40% and peripheral collisions 40–80% at energies $\sqrt{\mathrm{s}_{{}_{\mathrm{NN}}}}$leq 19.6 GeV is relatively lower as compared to central collisions and start to increase at energies at energies $\sqrt{\mathrm{s}_{{}_{\mathrm{NN}}}}$geq 27 GeV. On the other hand, $\langle\beta\rangle$ is slightly larger at the same centrality classes and energies. This may be related to the change in the collision geometry and different production mechanisms of light (anti-)nuclei at lower energies specifically in peripheral collisions where spectators plays an important role. Thus, a clear hierarchical structure seems to be seen in kinetic freeze-out suggesting more massive particles leave the medium earlier in time compared to those of light particles. It is important to note that the published data from the STAR experiment at RHIC on Au+Au collisions at $\sqrt{\mathrm{s}_{{}_{\mathrm{NN}}}}$ = 62.4 GeV and 200 GeV [73, 74, 75] used a slightly different configuration of the detector setup as compared to $\sqrt{\mathrm{s}_{{}_{\mathrm{NN}}}}$ = 7.7 – 39 GeV data sets. This can be seen as a slight difference between our estimations and the published results. The foreseen analysis of identified hadrons with recent detector upgrades will be helpful to understand this slight difference. We have also repeated the analysis by treating $n$ as a free parameter. It is shown that the best fit values of the extracted parameters $T_{kin}$ and $\langle\beta\rangle$ are found to be nearly unaltered when letting the parameter $n$ free for all energies.
@@
In the coalescence model nuclei are assumed to be formed at a later stage of the collisions precisely at or after the kinetic freeze-out. Under this condition, one can anticipate that the $p_{\rm T}$ spectra of both protons and (anti-)nuclei reflect a common temperature and velocity field of the source at or later stage of the nuclear cluster formation. To provide a more detailed explanation, we have categorized various particle species into distinct sets and performed the BW fit. Set A includes $\pi^{\pm}$, $K^{\pm}$, $p(\bar{p})$, $d(\bar{d})$ and $t$; Set B comprises $p(\bar{p})$, $d(\bar{d})$ and $t$; Set C consists of $d(\bar{d})$ and $t$; and Set D contains only $\pi^{\pm}$, $K^{\pm}$, $p(\bar{p})$. Note that we use ${}^{3}He$ instead of $t$ where no data for triton is available at LHC energies. The extracted parameters from different set of combinations in (0–10%) central Au+Au collisions at $\sqrt{\mathrm{s}_{{}_{\mathrm{NN}}}}$ = 27 GeV are listed in Table 2 as an example. The similar trend is observed for the rest of energies. It is evident from the Table 2 that including $p(\bar{p})$ in the fit with light (anti-)nuclei exhibits a similar velocity profile as observed when $\pi^{\pm}$, $K^{\pm}$, $p(\bar{p})$ are considered in the fit (Set A). However, there is a notable difference in the kinetic freeze-out temperature ($T_{kin}$) and is found to be significantly larger ($129\pm 9$) MeV which is unexpected in the context of final-state coalescence and requires further investigation. However, the observed results aligns with the hypothesis of statistical hadronization in which light (anti-)nuclei are formed close to the phase boundary without undergoing significant interactions at later stages of the system evolution. The observed scenario might be explained by considering prehadronic multiquark states and has been discussed in detail in Ref. [36, 79]. We repeated the exercise for Pb+Pb collisions at $\sqrt{\mathrm{s}_{{}_{\mathrm{NN}}}}$ = 2.76 TeV and observed similar findings. The corresponding extracted parameters are listed in Table 2. The velocity profile of Set B, when including $p(\bar{p})$ in the fit along with light (anti-)nuclei, appears to be relatively lower compared to Set A. However, $T_{kin}$ is significantly larger ($149\pm 7$) MeV. To understand this, we introduced another set, denoted as Set B, in which we keep the same velocity profile as that of Set A. As a result, the $T_{kin}$ is noticeably reduced to ($132\pm 5$) MeV and is consistent with ALICE experimental results from Pb+Pb collisions at $\sqrt{\mathrm{s}_{{}_{\mathrm{NN}}}}$ = 5.02 TeV [79].
@@ IV Conclusions
The analysis of transverse momentum $p_{\rm T}$ spectra of produced particles provide insights into the particle production mechanisms as well as the freeze-out conditions of the system. We reported the comprehensive investigation of the kinetic freeze-out properties of identified hadrons ($\pi^{\pm}$, $K^{\pm}$ and $p(\bar{p})$) along with light (anti-)nuclei $d(\bar{d})$, $t(\bar{t})$ and ${}^{3}He$ in relativistic heavy-ion collisions at RHIC and LHC energies. The $p_{\rm T}$ distribution of identified hadrons including light (anti-)nuclei is fitted simultaneously with the Blast-Wave model. This detailed study suggests that relatively heavier particles like light (anti-)nuclei participate in a common flow field resulting in a slight increase in $T_{kin}$ and slight decrease in $\langle\beta\rangle$ in comparison to those when only identified hadrons are included in the fit. This indicates that heavier particles decouple earlier in time from the fireball compared to the light hadrons due to their higher masses and are more sensitive to the effects of radial flow. The strong anti-correlation between $T_{kin}$ and $\langle\beta\rangle$, i.e. the value of $T_{kin}$ increases from central to peripheral collisions indicating shorter-lived fireball in peripheral collisions, The value of $\langle\beta\rangle$ decrease from central to peripheral collisions is consistent with the picture of more rapid expansion of the system in central collisions at both, RHIC and LHC energies. The extracted $T_{kin}$ show weak energy dependence at RHIC energies while decreases at the LHC energies and $\langle\beta\rangle$ shows increasing trend from RHIC to LHC energies.

### v3

@@ Abstract
We report the investigation of the kinetic freeze-out properties of identified hadrons ($\pi^{\pm}$, $K^{\pm}$ and $p(\bar{p})$) along with light (anti-)nuclei $d(\bar{d})$, $t(\bar{t})$ and ${}^{3}He$ in relativistic heavy-ion collisions at RHIC and LHC energies. A simultaneous fit is performed with the Blast-Wave (BW) model to the transverse momentum ($p_{\rm T}$) spectra of identified hadrons together with light (anti-)nuclei produced in Au+Au collisions at $\sqrt{\mathrm{s}_{{}_{\mathrm{NN}}}}$ = 7.7 – 200 GeV at the RHIC and in Pb+Pb collisions at $\sqrt{\mathrm{s}_{{}_{\mathrm{NN}}}}$ = 2.76 TeV at the LHC. The energy and centrality dependence of freeze-out parameters, i.e., kinetic freeze-out temperature ($T_{kin}$) and collective flow velocity $\langle\beta\rangle$ has been studied. It is observed that light (anti-)nuclei also participate in the collective expansion of the medium created in the collision when included in a common fit with the light hadrons. We observe a marginal rise in $T_{kin}$ and a slight decrease in $\langle\beta\rangle$ when compared to the values obtained from the fit to light hadrons. A similar $\langle\beta\rangle$ and significantly larger $T_{kin}$ is observed when a fit is performed to only protons and light (anti-)nuclei. Both, $T_{kin}$ and $\langle\beta\rangle$ show a weak energy dependence at most collision energies.
@@ III.1 Transverse Momentum ($p_{\rm T}$) Spectra
For light hadrons, we repeated the analysis presented in Ref. [56] for a simultaneous fit to the $p_{\rm T}$ spectra with BW model. The obtained parameters from the BW fit are consistent with the previously reported results [56]. The low $p_{\rm T}$ part of the pion spectra is significantly influenced by resonance decays therefore, $p_{\rm T}$ $<0.5$ GeV/$c$ is excluded from the fit. Conversely, at higher $p_{\rm T}$, eventual hard contributions may become significant, which are not anticipated to be effectively described by the BW model [65]. Therefore, the sensitivity of the BW model parameters especially $T_{kin}$ is contingent upon the choice of $p_{\rm T}$ ranges used for fitting [57]. The fit range for light hadrons have been limited according to Ref. [56] (0.5 – 1.3 GeV/$c$, 0.25 – 1.4 GeV/$c$, 0.4 – 1.3 GeV/$c$ for $\pi^{\pm},K^{\pm}$ and $p(\bar{p})$ respectively). However, for light (anti-)nuclei the spectrum is fitted up to the maximum available $p_{\rm T}$ range. The fit ranges can significantly affect the extracted $T_{kin}$, especially at lower $p_{\rm T}$. Further details on fit ranges are reported in Refs. [70, 71].
@@ III.2 Kinetic Freeze-out Parameters
Figures 3 and  4 show the centrality dependence of temperature ($T_{kin}$) and radial flow velocity ($\langle\beta\rangle$), respectively. Parameters are extracted by simultaneous fits to the $p_{\rm T}$ spectra of light hadrons along with light (anti-)nuclei in Au+Au collisions at $\sqrt{\mathrm{s}_{{}_{\mathrm{NN}}}}$ = 7.7 – 200 GeV from RHIC [56] and Pb+Pb collisions at $\sqrt{\mathrm{s}_{{}_{\mathrm{NN}}}}$ = 2.76 TeV [57] from LHC. For simplicity, the centrality classes of $\pi^{\pm}$, $K^{\pm}$, $p(\bar{p})$ are merged and refitted to make it consistent with those used for $d(\bar{d})$ and $t(\bar{t})$ measurements, i.e. 0–10%, 10–20%, 20–40% and 40–80%. Due to unavailability of the triton $p_{\rm T}$ spectra from Pb+Pb collisions at $\sqrt{\mathrm{s}_{{}_{\mathrm{NN}}}}$ = 2.76 TeV, we included ${}^{3}He$ in the fits instead. Blue markers represent the published results for $\pi^{\pm}$, $K^{\pm}$, and $p(\bar{p})$. Green markers depict $\pi^{\pm}$, $K^{\pm}$, and $p(\bar{p})$ data with combined centrality classes. Red markers correspond to $T_{kin}$ of $\pi^{\pm}$, $K^{\pm}$, $p(\bar{p})$, and $d(\bar{d})$. It is evident from Fig. 3 that $T_{kin}$ exhibits an increasing trend as a function of centrality for all collision energies. Furthermore, the inclusion of heavier particle species, such as light (anti-)nuclei, results in slightly higher temperatures compared to those of light hadrons. This effect is more prominent at energies $\sqrt{\mathrm{s}_{{}_{\mathrm{NN}}}}$geq 62.4 GeV. Fig. 4 shows that fitting both light hadrons with light (anti-)nuclei together leads to a smaller values of $\langle\beta\rangle$ as compared to those when only light hadrons are considered in the fits. To study the influence of light (anti-)nuclei data on the global characteristics, we repeat the fit by simultaneously excluding $d(\bar{d})$ and $t$. This exhibits similar trend, but with slightly lower value of $T_{kin}$ and slightly larger values of $\langle\beta\rangle$. Other effects, like elliptic flow ($v_{2}$) of deutrons can also be used study the distinction between coalescence and statistical production. More details are discussed in Ref. [78]. Moreover, the $T_{kin}$ in mid-central 20–40% and peripheral collisions 40–80% at energies $\sqrt{\mathrm{s}_{{}_{\mathrm{NN}}}}$leq 19.6 GeV is relatively lower as compared to central collisions and start to increase at energies at energies $\sqrt{\mathrm{s}_{{}_{\mathrm{NN}}}}$geq 27 GeV. On the other hand, $\langle\beta\rangle$ is slightly larger at the same centrality classes and energies. This may be related to the change in the collision geometry and different production mechanisms of light (anti-)nuclei at lower energies specifically in peripheral collisions where spectators plays an important role. Thus, a clear hierarchical structure seems to be seen in kinetic freeze-out parameters. We have also repeated the analysis by treating $n$ as a free parameter. It is shown that the best fit values of the extracted parameters $T_{kin}$ and $\langle\beta\rangle$ are found to be nearly unaltered when letting the parameter $n$ free for all energies.
@@
In the coalescence model nuclei are assumed to be formed at a later stage of the collisions precisely at or after the kinetic freeze-out. Under this condition, one can anticipate that the $p_{\rm T}$ spectra of both protons and (anti-)nuclei reflect a common temperature and velocity field of the source at or later stage of the nuclear cluster formation. To provide a more detailed explanation, we have categorized various particle species into distinct sets and performed the BW fit. Set A includes $\pi^{\pm}$, $K^{\pm}$, $p(\bar{p})$, $d(\bar{d})$ and $t$; Set B comprises $p(\bar{p})$, $d(\bar{d})$ and $t$; Set C consists of $d(\bar{d})$ and $t$; and Set D contains only $\pi^{\pm}$, $K^{\pm}$, $p(\bar{p})$. Note that we use ${}^{3}He$ instead of $t$ where no data for triton is available at LHC energies. The extracted parameters from different set of combinations in (0–10%) central Au+Au collisions at $\sqrt{\mathrm{s}_{{}_{\mathrm{NN}}}}$ = 27 GeV are listed in Table 2 as an example. The similar trend is observed for the rest of energies. It is evident from the Table 2 that including $p(\bar{p})$ in the fit with light (anti-)nuclei exhibits a similar velocity profile as observed when $\pi^{\pm}$, $K^{\pm}$, $p(\bar{p})$ are considered in the fit (Set A). However, there is a notable difference in the kinetic freeze-out temperature ($T_{kin}$) and is found to be significantly larger ($129\pm 9$) MeV which is unexpected in the context of final-state coalescence and requires further investigation. The observed results aligns with the hypothesis of statistical hadronization in which light (anti-)nuclei are formed close to the phase boundary without undergoing significant interactions at later stages of the system evolution. The observed scenario might be explained by considering pre-hadronic multiquark states and has been discussed in detail in Ref. [36, 79]. We repeated the exercise for Pb+Pb collisions at $\sqrt{\mathrm{s}_{{}_{\mathrm{NN}}}}$ = 2.76 TeV and observed similar findings. The corresponding extracted parameters are listed in Table 2. The velocity profile of Set B, when including $p(\bar{p})$ in the fit along with light (anti-)nuclei, appears to be relatively lower compared to Set A. However, $T_{kin}$ is significantly larger ($149\pm 7$) MeV. To understand this, we introduced another set, denoted as Set B, in which we keep the same velocity profile as that of Set A. As a result, the $T_{kin}$ is noticeably reduced to ($132\pm 5$) MeV and is consistent with ALICE experimental results from Pb+Pb collisions at $\sqrt{\mathrm{s}_{{}_{\mathrm{NN}}}}$ = 5.02 TeV [79].
@@ IV Conclusions
The analysis of transverse momentum $p_{\rm T}$ spectra of produced particles provide insights into the particle production mechanisms as well as the freeze-out conditions of the system. We reported the comprehensive investigation of the kinetic freeze-out properties of identified hadrons ($\pi^{\pm}$, $K^{\pm}$ and $p(\bar{p})$) along with light (anti-)nuclei $d(\bar{d})$, $t(\bar{t})$ and ${}^{3}He$ in relativistic heavy-ion collisions at RHIC and LHC energies. Blast-Wave analysis of the $p_{\rm T}$ distribution of identified hadrons together with light (anti-)nuclei results in a slight increase in $T_{kin}$ and slight decrease in $\langle\beta\rangle$ in comparison to those when only identified hadrons are included in the fit. This suggests that relatively heavier particles like light (anti-)nuclei participate in a common flow field. Including only $p(\overline{p})$ in the fit together with light (anti-)nuclei leads to a surprisingly large extracted $T_{kin}$. The strong anti-correlation between $T_{kin}$ and $\langle\beta\rangle$, i.e., the value of $T_{kin}$ increases from central to peripheral collisions, indicating a shorter-lived fireball in peripheral collisions. Conversely, the value of $\langle\beta\rangle$ decreases from central to peripheral collisions, consistent with the picture of a more rapid expansion of the system in central collisions at both RHIC and LHC energies. At RHIC energies, the extracted $T_{kin}$ exhibits weak energy dependence. However, it decreases at LHC energies, likely due to the longer scattering phase at higher energies. On the other hand, $\langle\beta\rangle$ shows an increasing trend from RHIC to LHC energies. Further investigations into other important aspects, such as flow, can be used to distinguish between different production mechanisms of light (anti-)nuclei.